how do you graph x^2+y^2+12x-6y-9=0
Answer by lwsshak3(11628)

how do you graph x^2+y^2+12x-6y-9=0
..
x^2+y^2+12x-6y-9=0
completing the square
(x^2+12x+36)+(y^2-6y+9)=0+36+9=45
(x+6)^2+(y-9)^2=45
This equation is a circle with center at (-6,9) and radius=√45=6.7..
see graph below:
..
y=(45-(x+6)^2)^.5+9